Opened 8 years ago

Closed 8 years ago

Last modified 3 years ago

#362 closed flight processing (fixed)

EUFAR10/07 (ADDRESSS), flight day 234d, Summer School Geo Site

Reported by: iopa Owned by: knpa
Priority: alpha 4 high Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by iopa)

Data location: ~arsf/workspace/EUFAR10_07-2010_234d_Summer_School_Geo_Site

Data arrived from ARSF via SATA disk Q on 15/09/2010.

Scientific objective: Mapping of lithology, fault strucuture,
erosion parameters and soil components

Priority: a4h

PI: A. Zlinszky

Note: Data is currently symlinked from ~arsf/arsf_data/2010/flight_data/archive_overflow and when this is resolved the symlinks will need to be re-directed.

PI has requested that data be made available via NEODC promptly as well as normal delivery - data will be downloaded from NEODC by students.

Sensors:

  • Eagle (04/03/2011)
  • Hawk (04/03/2011)
  • Leica LIDAR (28/03/2011)
  • Digital Photography (04/03/2011)

Change History (32)

comment:1 Changed 8 years ago by iopa

  • Description modified (diff)

comment:2 Changed 8 years ago by knpa

  • Description modified (diff)

Added scientific objective.

comment:3 Changed 8 years ago by knpa

Appears to be an issue with the bandset labels in the header files; they start at 900nm which is lower than it should be. Possibly incorrect bandset selected or none specified. Will leave this until end of the year and new Hawk calibration.

comment:4 Changed 8 years ago by benj

  • Description modified (diff)

comment:5 Changed 8 years ago by benj

  • Description modified (diff)

PI has requested that data be made available via NEODC promptly as well as normal delivery - data will be downloaded from NEODC by students.

comment:6 Changed 8 years ago by emca

Processed the camera imagery from raw to tiff format

comment:7 Changed 8 years ago by anch

Applanix has been processed for 238e which covers this flight, so proc has been copied into the applanix folder from 238e

comment:8 Changed 8 years ago by anch

Starting hyperspectral processing

comment:9 Changed 8 years ago by anch

Submitted hyperspectral for delivery check

comment:10 Changed 8 years ago by anch

Starting LiDAR processing.

comment:11 Changed 8 years ago by anch

IPAS from 234e covers this flight so I have copied it from there.

comment:12 Changed 8 years ago by iopa

Checking hyperspectral.

comment:13 Changed 8 years ago by iopa

Hyperspectal checking done;

  • there's been a keyboard slip somewhere, proj code is misspelt in delivery directory, DEM name, logsheet, and DEM name within the readme.
  • eagle/hawk flightline id's in readme for line 3 should be 012

else the data seems good and ready to deliver.

comment:14 Changed 8 years ago by emca

Starting camera processing

comment:15 Changed 8 years ago by anch

Submitted lidar for delivery check

comment:16 Changed 8 years ago by emca

Camera ready for delivery check at
~airborne/workspace/EUFAR10_07-2010_234d_Summer_School_Geo_Site/delivery/20110303

comment:17 Changed 8 years ago by anch

Checking camera

comment:18 Changed 8 years ago by anch

Camera delivery is fine and ready to deliver

comment:19 Changed 8 years ago by knpa

  • Owner set to anch
  • Status changed from new to assigned

Taking lidar delivery check.

comment:20 Changed 8 years ago by anch

  • Description modified (diff)

Hyperspectral, and camera sent to PI

comment:21 Changed 8 years ago by knpa

LIDAR delivery check complete.

All looks fine apart from:

  1. lidar ascii files have 'e' in name rather than 'd'. Assuming this is just a typo and these files are for the right project :) Also needs to be fixed in Readme contents section.
  2. EUFARA10_07-2010_234d_Summer_School_Geo_Site.pdf - project code spelt wrong

comment:22 Changed 8 years ago by anch

  • Summary changed from EUFAR10/07, flight day 234d, Summer School Geo Site to EUFAR10/07 (ADDRESSS), flight day 234d, Summer School Geo Site

comment:23 Changed 8 years ago by iopa

  • Description modified (diff)

LiDAR has been dispatched (28/03/2011)

comment:24 Changed 8 years ago by emca

  • Component changed from Processing: general to Archiving

All data has been rsync'd to
/users/rsg/arsf/arsf_data/2010/flight_data/archive-overflow/EUFAR10_07-2010_234d_Summer_School_Geo_Site

Ready for archiving

comment:25 Changed 8 years ago by emca

LAS files are included with the delivery folder, so have been removed from leica/proc_laser to avoid duplication.

This project had aimms, grimm and tabi data.


comment:26 Changed 8 years ago by emca

Archive tarball created and made available to NEODC

comment:27 Changed 8 years ago by knpa

  • Owner changed from anch to knpa

Backed up by NEODC, moved to archive disk.

comment:28 Changed 8 years ago by knpa

  • Milestone set to 2010 data processing completion
  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace version. Closing ticket.

comment:29 Changed 7 years ago by adbe

Hyperspectral Delivery Check:

  • Mention dropped scans in ReadMe
  • Hawk 3,5,6 mapped filesize does not match with header

comment:30 Changed 6 years ago by anhi

Hawk lines 3,5,6 remade and put into delivery

Last edited 6 years ago by anhi (previous) (diff)

comment:31 Changed 6 years ago by anhi

Reprocessing hyperspectral delivered to Andras Zlinszky on 12/09/12 on disk 131.

Note: See TracTickets for help on using tickets.